\newpage

\begin{thebibliography}{99}

\bibitem{malda}
J. M. Maldacena:
''The Large N Limit of Superconformal Field Theories and Supergravity''
Adv. Theor. Math. Phys. {\bf 2} (1998) 231.

\bibitem{witten}
E. Witten:''Anti de Sitter space and holography''. Adv. Theor. Math. Phys. {\bf 2} (1998) 253

\bibitem{gubser}
S. S. Gubser, I.R. Klebanov and A.M. Polyakov: ''Gauge theory correlators from
non-critical string theory''. Phys. Lett. B {\bf 428} (1998) 105.

\bibitem{dani} U.H. Danielsson, E. Keski-Vakkuri and M. Kruczenski," Vacua, propagators, and
holographic probes in AdS/CFT". J. High Energy Phys. 9901 (1999) 002.

\bibitem{ry} S. Ryang, The Hadamard function and the Feynman propagator in the AdS/CFT
correspondence. Phys. Lett. B {\bf } 469 (1999) 87. 

\bibitem{kle} I. R. Klebanov and E. Witten. “AdS/CFT Correspondence and Symmetry
Breaking”. Nuc. Phys. B {\bf 556} (1999) 89.

\bibitem{pa} Pablo Minces and Victor O. Rivelles. “Scalar Field Theory in the AdS/CFT Correspondence
Revisited”. Nuc. Phys. B {\bf 572} (2000) 651.

\bibitem{wo} Wolfgang Mück. “Studies on the adS/CFT correspondence”. PhD thesis. Simon Fraser
University, 1999.
\url{http://www.collectionscanada.gc.ca/obj/s4/f2/dsk1/tape9/PQDD_0025/NQ51904.pdf}

\bibitem{wo1}  Wolfgang Mück and K. S. Viswanathan. “Conformal Field Theory Correlators from
Classical Scalar Field Theory on $AdS_{d + 1}$”
Phys. Rev D {\bf 58} (1998) 041901(R).

\bibitem{er} H. Erbin:"Scalar propagators on adS space"
\url{https://www.lpthe.jussieu.fr/~erbin/files/ads_propagators.pdf}

\bibitem{bala1}
V. Balasubramanian, P. Kraus, A. Lawrence and S.P. Trivedi: ''Holographic probes of
anti-de Sitter spacetimes''. Phys. Rev. D {\bf 59} (1999) 104021.

\bibitem{bala2}
V. Balasubramanian, S.B. Giddings and A. Lawrence:''What do CFTs tell us about
anti-de Sitter spacetimes?''. JHEP  {\bf 9903} (1999) 001.

\bibitem{bala3}
V. Balasubramanian, P. Kraus and A. Lawrence:''Bulk versus boundary dynamics in
anti-de Sitter spacetime''. Phys. Rev. D {\bf 59} (1999) 046003.

\bibitem{son} 
D. T. Son and A. O. Starinets:''Minkowski-space correlators in AdS/CFT
correspondence: recipe and applications''. JHEP {\bf 09} (2002) 042.

\bibitem{tp4} L. Schwartz: ''Th\'eorie des distributions''.
Hermann, Paris (1966)

\bibitem{tp7} I. M. Gel'fand and G. E. Shilov: ''Generalized
Functions'' {\bf Vol. 1}. Academic Press (1964).

\bibitem{tp17} S. Bochner: ''Lectures on Fourier Integrals''.Princeton University
Press, Princeton, NJ, 1959.

\bibitem{tp13} I. S. Gradshteyn and I. M. Ryzhik:''Table of Integrals,
Series and Products''. Academic Press, Inc (1980).

\bibitem{jones} D. S. Jones:''Generalised Functions''. Cambridge University Press
(1966).

\bibitem{nas}
H. Nastase:''Introduction to the AdS/CFT Correspondence''.
Cambridge University Press (2015).

\bibitem{z2} D. Z. Freedman, S. D. Mathur, A. Matusis and L. Rastelli.
Nuc. Phys. B {\bf 546}, 96 (1999).

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{graphics}
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amstext}
\usepackage{amscd}
\usepackage{amsfonts}
\usepackage{appendix}
\usepackage{hyperref}
\usepackage{textcomp}
\usepackage{color}
\DeclareMathAlphabet{\EuFrak}{U}{euf}{m}{n}
\DeclareMathAlphabet{\EuScript}{U}{eus}{m}{n}

\newcommand{\nd}{\noindent}

\title{{\bf  A rigorous calculation of the Feynman and Wheeler Scalar Fields propagators in the ADS / CFT 
correspondence using distribution theory}}
\author{{A. Plastino$^{1,3,4}$, M.C.Rocca$^{1,2,3}$}\\
\small{$^1$ Departamento de F\'{\i}sica,
Universidad Nacional de La Plata,}\\
\small{$^2$ Departamento de Matem\'{a}tica,
Universidad Nacional de La Plata,}\\
\small{$^3$ Consejo Nacional de Investigaciones Cient\'{\i}ficas
y Tecnol\'{o}gicas}\\
\small{(IFLP-CCT-CONICET)-C. C. 727, 1900 La Plata -
Argentina}\\\small{$^4$  SThAR - EPFL, Lausanne, Switzerland}}

\date{\today}

\begin{document}

\maketitle

\vspace{-5mm}

\begin{abstract}
By appeal to Distribution Theory  we discuss in rigorous fashion,   
without appealing to {\bf any conjecture} (as usually done by other authors),
the boundary-bulk propagators for the scalar field, both in the non-massive  and
massive cases.  These calculations, new in the literature as far as we know, are carried out in  two instances: (i)  
when the boundary is a Euclidean space and (ii) when it is of Minkowskian nature. 
In this last case we compute also  three propagators: Feynman's, Anti-Feynman's, and Wheeler's 
(half advanced plus half retarded). For an operator corresponding to scalar field we
obtain the two points correlations functions in the three instances above mentioned\\
\nd {\bf PACS}11.25.-w, 04.60.Cf, 02.30.Sa, 03.65.Db.\\
\nd {\bf KEYWORDS} Distribution theory, ADS/CFT correspondence; Boundary-bulk propagators;
Feynman's propagators, Wheeler's propagators.\\

\end{abstract}

\newpage

\tableofcontents

\newpage

\renewcommand{\theequation}{\arabic{section}.\arabic{equation}}

\section{Introduction}

\nd
Propagators and correlators are one of the essential tools to work,
for example, in Quantum Field Theory (QFT) and String Theory (ST).
In particular, in formulating the correspondence ADS/CFT (Anti-de Sitter/
Conformal Field Theory)/ This correspondence was established by
Maldacena \cite{malda} in 1998 and is universally regarded as a very useful model
for many purposes.

\nd
The bibliography on this subject, for scalar fields, is quite extensive.
We give here just a small part of it in \cite{witten,gubser,dani,ry,kle,pa,wo,wo1}.
For a more complete bibliography the reader is directed to the report \cite{er}
\nd
One of the ADS/CFT correspondence's prescriptions (see \cite{witten}) will 
allow us to evaluate the correlators on the boundary of ADS space 
The first boundary-bulk propagator was calculated by Witten a few months after the appearance of \cite{malda}, entitled  {\it Anti de Sitter space and holography}. In this case the boundary is an Euclidean space  
\cite{witten,gubser}.

\nd 
By appeal to distribution theory, in this work  we evaluate instead  the boundary-bulk propagators
for the case in which the boundary is a Minkowskian space. Distribution theory is not a tool employed by AdS/CFT practitioners, as far as we know.

\nd
In such regard, remark  that some attempts have been made before in \cite{bala1,bala2,bala3}.
The only previous (and almost ) attempt to try to calculate boundary-bulk
propagators in the Minkowskian boundary for the Anti-de Sitter space [in the ADS/CFT correspondence] was made by  Son and Starinets 
(SS) in  2002  \cite{son}. However, 
 SS needed to formulate   a conjecture that we show here to become unnecessary if one uses the full
distributions-theory of type $S^{'}$ (of Schwartz).
This important  work was entitled ''Minkowski-space correlators in AdS/CFT
correspondence: recipe and applications''.   We must also mention the work of Freedman et al. \cite{z2}, in which the authors
deal with the case of a Euclidean boundary.  
Freedman, however,  did not treat the case of a Minkowskian boundary,
at least in the way that Son and Starinets did.
Note that we make full use of distribution theory. This does not entail, of course, a simple $i\epsilon$ prescription, but a much more elaborate treatment, that has not been performed before in this field.
Let us remark, as this is an important point for us, that in this paper we do not evaluate renormalized correlation functions. 
We will do that in a forthcoming  paper using the method given in \cite{z2}. 

\nd
Thus, in the present effort we evaluate, without any conjecture, the boundary-bulk
propagators corresponding to the following  three cases i) Feynman, ii) Anti-Feynman, and iii)  Wheeler.
 We do this both, for a massless scenario  and for  the massive ones,
(a scalar field involved). 
Later we calculate the two points correlators (TPC)  for operators corresponding to this scalar 
field in the three instances previously mentioned.
We clarify that in this paper we do not evaluate the renormalized TPC.
We will do that in a next paper using the method given in \cite{z2}. 

\nd In the present effort  we demonstrate that the Feynman propagator must be a function of $\rho+i0$ (see below for the notation) in  momentum space, and therefore a function of $x^2-i0$ in configuration space. We show that something similar happens with the Anti-Feynman propagator.  {\it For the first time ever}, we calculate  the Wheeler's propagator as well.
{\it Note that, until the 90's, the only field propagators that had been  calculated were 
Anti-de Sitter (spatial) ones.}

\nd
The paper is organized as follows: Section 2 deals with with the Euclidean case.
In it, the three different propagators referred to above can not be distinguished
(neither  in the massive nor in the massless instances).

\nd
In section 3 we tackle similar scenarios as those of section 2, but now
in Minkowski's space, where the three propagators can be distinguished.

\nd
In section 4 we compute in Euclidean space the TPC for a scalar operator corresponding to a scalar field via
Witten's prescription

\nd
In section 5 we generalize the calculations of section 4 to Minkowski's
space. We obtain in this fashion the two-points correlations functions
corresponding to the three different propagators of our list above.

\nd
Finally, some conclusions are drawn in section 6.   

\section{Euclidean Case}

\subsection{Massless Scalar Field Propagator}

\setcounter{equation}{0}

The Klein-Gordon equation in $ADS_{\nu+1}$
for the scalar field $\phi(z,\vec{x})$ reads,  in Poincare coordinates,
\begin{equation}\label{ep2.1}
z^2\partial_z^2\phi(z,\vec{x})+(1-\nu)z\partial_z\phi(z,\vec{x})+z^2\nabla^2\phi(z,\vec{x})-
\Delta(\Delta-\nu)\phi(z,\vec{x})=0,
\end{equation}
where $\Delta(\Delta-\nu)\geq 0$ plays the role of $m^2$. We exclude tachyons from of this
treatment.
Here $\Delta$ is the conformal dimension, $\nu$ the boundary's dimension,
and $\vec {x}$ their coordinates.
The Fourier transform in the 
variables $\vec{x}$ of the field $\phi(z,\vec{x})$ is
\begin{equation}
\label{ep2.2}
\hat{\phi}(z,\vec{k})=\int \phi(z,\vec{x}) e^{i\vec{k}\cdot \vec{x}} d^{\nu}x. 
\end{equation}
Using (\ref{ep2.2}),  (\ref{ep2.1}) takes the form
\begin{equation}
\label{ep2.3} 
z^2\partial_z^2\hat{\phi}(z,\vec{k})+(1-\nu)z\partial_z\hat{\phi}(z,\vec{k})-[z^2k^2+
\Delta(\Delta-\nu)]\hat{\phi}(z,\vec{k})=0.
\end{equation}

We analyze now the massless case given by $\Delta=0, \nu$.
For it we have  the motion equation
\begin{equation}
\label{ep2.4}
z^2\partial_z^2\hat{\phi}(z,\vec{k})+(1-\nu)z\partial_z\hat{\phi}(z,\vec{k})-z^2k^2\hat{\phi}(z,\vec{k})=0,
\end{equation}
or equivalently (for $z\neq 0$)
\begin{equation}
\label{ep2.5}
\partial_z^2\hat{\phi}(z,\vec{k})+\frac {(1-\nu)} {z}\partial_z\hat{\phi}(z,\vec{k})-k^2\hat{\phi}(z,\vec{k})=0.
\end{equation}
In the variable $ z $,  this equation is of the Bessel type (see \cite{tp13})
\begin{equation}
\label{ep2.6}
F''(z)+\frac {(1-2\alpha)} {z}F'(z)-\left[k^2+\frac {\mu^2-\alpha^2} {z^2}\right]F(z)=0
\end{equation}
The pertinent solution (that does not diverge when the argument tends to infinity) is 
\begin{equation}
\label{ep2.7}
F(z)=z^{\alpha}{\cal K}_\mu(kz).
\end{equation}
Thus, the solution of (\ref{ep2.5}) becomes
\begin{equation}
\label{ep2.8}
{\hat{\phi}}(z,k)=z^{\frac {\nu} {2}}{\cal K}_{\frac {\nu} {2}}(kz).
\end{equation}
One easily verifies that, for infinitesimal $z$  \cite{tp13},
\begin{equation}
\label{ep2.9}
{\cal K}_{\frac {\nu} {2}}(kz)=\frac {2^{\frac {\nu} {2}-1}\Gamma\left(\frac {\nu} {2}\right)}
{(kz)^{\frac {\nu} {2}}}+O\left((kz)^{-\frac {\nu} {2}+2}\right),
\end{equation}
and therefore
\begin{equation}
\label{ep2.10}
\lim_{z\rightarrow 0}z^{\frac {\nu} {2}}{\cal K}_{\frac {\nu} {2}}(kz)=
\frac {2^{\frac {\nu} {2}-1}\Gamma\left(\frac {\nu} {2}\right)}
{k^{\frac {\nu} {2}}}.
\end{equation}
In other words, the solution is regular at the origin and vanishes at infinity
(in the variable $z$). Accordingly,  we have, for the field in the bulk, the solution
\begin{equation}
\label{ep2.11}
\phi(z,\vec{x})=\frac {z^{\frac {\nu} {2}}} {(2\pi)^{\nu}}
\int a(\vec{k}) {\cal K}_{\frac {\nu} {2}}(kz)
e^{-i\vec{k}\cdot \vec{x}}d^{\nu}k.
\end{equation}
This solution must  reduce itself to the field $\phi_0(x)$ on the boundary, so that
%%Thus we have
\[\phi(0,\vec{x})=\phi_0(\vec{x})=\frac {2^{\frac {\nu} {2}-1}
\Gamma\left(\frac {\nu} {2}\right)} {(2\pi)^{\nu}}
\int a(\vec{k}) k^{-\frac {\nu} {2}}
e^{-i\vec{k}\cdot \vec{x}}d^{\nu}k=\]
\begin{equation}
\label{ep2.12}
\frac {1} {(2\pi)^\nu}
\int \hat{\phi}_0(\vec{k})e^{-i\vec{k}\cdot \vec{x}}d^{\nu}k.
\end{equation}
From this last equation we can obtain $a(k)$ as a function of $\hat{\phi}_0$ and
then write
\begin{equation}
\label{ep2.13}
\phi(z,\vec{x})=\frac {z^{\frac {\nu} {2}}2^{1-\frac {\nu} {2}}} 
{(2\pi)^{\nu}\Gamma{\left(\frac {\nu} {2}\right)}}
\int k^{\frac {\nu} {2}} {\cal K}_{\frac {\nu} {2}}(kz)
\hat{\phi}_0(\vec{k})
e^{-i\vec{k}\cdot \vec{x}}d^{\nu}k,
\end{equation}
or, equivalently,
\begin{equation}
\label{ep2.14}
\phi(z,\vec{x})=\frac {z^{\frac {\nu} {2}}2^{1-\frac {\nu} {2}}} 
{(2\pi)^{\nu}\Gamma{\left(\frac {\nu} {2}\right)}}
\iint k^{\frac {\nu} {2}} {\cal K}_{\frac {\nu} {2}}(kz)
{\phi}_0({\vec{x}}^{'})
e^{-i\vec{k}\cdot (\vec{x}-{\vec{x}}^{'})}d^{\nu}kd^{\nu}x^{'}.
\end{equation}
From (\ref{ep2.14}) we then obtain an expression of the boundary-bulk propagator 

\begin{equation}
\label{ep2.15}
K(z,\vec{x}-\vec{x}^{'})=
\frac {z^{\frac {\nu} {2}}2^{1-\frac {\nu} {2}}} 
{(2\pi)^{\nu}\Gamma{\left(\frac {\nu} {2}\right)}}
\int  k^{\frac {\nu} {2}} {\cal K}_{\frac {\nu} {2}}(kz)
e^{-i\vec{k}\cdot (\vec{x}-{\vec{x}}^{'})}d^{\nu}k.
\end{equation}
To carry out the integration in the variable $k$ we appeal to
the expressions for the  Fourier transform and its inverse obtained by
Bochner \cite{tp17}. For the Fourier transform we have
\begin{equation}
\label{ep2.16}
\hat{f}(k)=\int f(\vec{x})e^{i\vec{k}\cdot\vec{x}}d^{\nu}x=
\frac {(2\pi)^{\frac {\nu} {2}}} {k^{\frac {\nu} {2}-1}}
\int\limits_0^{\infty} r^{\frac {\nu} {2}} {\cal J}_{\frac {\nu} {2}-1}(kr)f(r) dr,
\end{equation}
and for its inverse
\begin{equation}
\label{ep2.17}
f(r)=\frac {1} {(2\pi)^{\nu}}
\int \hat{f}(\vec{k})e^{-i\vec{k}\cdot\vec{x}}d^{\nu}k=
\frac {1} {(2\pi)^{\frac {\nu} {2}}r^{\frac {\nu} {2}-1}}
\int\limits_0^{\infty} k^{\frac {\nu} {2}} {\cal J}_{\frac {\nu} {2}-1}(kr)\hat{f}(k) dk.
\end{equation}
Using these  relations   we have now 

\begin{equation}
\label{ep2.18}
K(z,\vec{x}-\vec{x}^{'})=
\frac {z^{\frac {\nu} {2}}2^{1-\frac {\nu} {2}}} 
{(2\pi)^{\nu}\Gamma{\left(\frac {\nu} {2}\right)}}
\frac {(2\pi)^{\frac {\nu} {2}}} {|\vec{x}-\vec{x}^{'}|^{\frac {\nu} {2}-1}} 
\int\limits_0^{\infty} k^\nu {\cal K}_{\frac {\nu} {2}}(kz)
{\cal J}_{\frac {\nu} {2}-1}(k|\vec{x}-\vec{x}^{'}|)dk.
\end{equation}
So as to evaluate the last integral we appeal  to a result from \cite{tp13}
\begin{equation}
\label{ep2.19}
\int\limits_0^{\infty} x^{\mu+\nu+1} {\cal K}_{\nu}(bx)
{\cal J}_{\mu}(ax)dx=2^{\mu+\nu}a^{\mu}b^{\nu}
\frac {\Gamma(\mu+\nu+1)} {(a^2+b^2)^{\mu+\nu+1}},
\end{equation}
Our deduction follows a different, simpler and complete path
than that of \cite{witten}. Our approach also has a didactic utility.
\begin{equation}
\label{ep2.20}
K(z,\vec{x}-\vec{x}^{'})=\frac {\Gamma(\nu)} {\pi^{\frac {\nu} {2}}
\Gamma\left(\frac {\nu} {2}\right)}
\left[\frac {z} {z^2+(\vec{x}-\vec{x}^{'})^2}\right]^\nu,
\end{equation}
which leads to 

\begin{equation}
\label{ep2.21}
\phi(z,\vec{x})=
\int K(z,\vec{x}-{\vec{x}}^{'})
{\phi}_0({\vec{x}}^{'})d^{\nu}x^{'},
\end{equation}
an  expression that, in turn, leads to
\begin{equation}
\label{ep2.22}
\lim_{z\rightarrow 0} K(z,\vec{x}-{\vec{x}}^{'})=
\delta(\vec{x}-{\vec{x}}^{'}).
\end{equation}

\subsection{Massive Field Propagator}

We now consider the massive case $\Delta\neq 0,\nu$.
The equation of motion for this case reads
\begin{equation}
\label{ep2.23}
z^2\partial_z^2\hat{\phi}(z,\vec{k})+(1-\nu)z\partial_z\hat{\phi}(z,\vec{k})-
[z^2k^2+\Delta(\Delta-\nu)]\hat{\phi}(z,\vec{k})=0,
\end{equation}
or equivalently,
\begin{equation}
\label{ep2.24}
\partial_z^2\hat{\phi}(z,\vec{k})+\frac {(1-\nu)} {z}\partial_z\hat{\phi}(z,\vec{k})-
\left[k^2+\frac {\Delta(\Delta-\nu)} {z^2}\right]\hat{\phi}(z,\vec{k})=0.
\end{equation}
The solution for this last equation is
\begin{equation}
\label{ep2.25}
{\hat{\phi}}(z,k)=z^{\frac {\nu} {2}}{\cal K}_\mu(kz),
\end{equation}
with
\begin{equation}
\label{ep2.26}
\mu=\pm\sqrt{\frac {\nu^2} {4}+\Delta(\Delta-\nu)}.
\end{equation}
Since  ${\cal K}_\mu(z)={\cal K}_{-\mu}(z)$, we select for $\mu$ in (\ref{ep2.26})
the plus sign. We have then
\begin{equation}
\label{ep2.27}
\phi(z,\vec{x})=\frac {z^{\frac {\nu} {2}}} {(2\pi)^{\nu}}
\int a(\vec{k}) {\cal K}_\mu(kz)
e^{-i\vec{k}\cdot \vec{x}}d^{\nu}k.
\end{equation}
For $\Delta\neq 0$, this solution is not regular at the origin.
To overcome this problem we select
\[\phi(\epsilon,\vec{x})=\phi_{\epsilon}(\vec{x})=
\frac {{\epsilon}^{\frac {\nu} {2}}} {(2\pi)^{\nu}}
\int a(\vec{k}) {\cal K}_\mu(k\epsilon)
e^{-i\vec{k}\cdot \vec{x}}d^{\nu}k=\]
\begin{equation}
\label{ep2.28}
\frac {1} {(2\pi)^\nu}
\int \hat{\phi}_{\epsilon}(\vec{k})e^{-i\vec{k}\cdot \vec{x}}d^{\nu}k,
\end{equation}
where $\epsilon$ is infinitesimal. From (\ref{ep2.28}) we have then
\begin{equation}
\label{ep2.29}
a(\vec{k})=\frac {\hat{\phi}_{\epsilon}(\vec{k})}
{{\epsilon}^{\frac {\nu} {2}}{\cal K}_\mu(k\epsilon)}.
\end{equation}
Replacing the result of (\ref{ep2.29}) into (\ref{ep2.27}) we obtain
\begin{equation}
\label{ep2.30}
\phi(z,\vec{x})=
\frac {1} {(2\pi)^{\nu}}
\left(\frac {z} {\epsilon}\right)^{\frac {\nu} {2}} 
\int \frac {{\cal K}_\mu(kz)}  {{\cal K}_\mu(k\epsilon)}
{\hat{\phi}_{\epsilon}(\vec{k})}
e^{-i\vec{k}\cdot \vec{x}}d^{\nu}k,
\end{equation}
or similarly,
\begin{equation}
\label{ep2.31}
\phi(z,\vec{x})=
\frac {1} {(2\pi)^{\nu}}
\left(\frac {z} {\epsilon}\right)^{\frac {\nu} {2}} 
\int\int \frac {{\cal K}_\mu(kz)}  {{\cal K}_\mu(k\epsilon)}
\phi_{\epsilon}(\vec{x}^{'})
e^{-i\vec{k}\cdot (\vec{x}-\vec{x}^{'})}d^{\nu}kd^{\nu}x^{'}.
\end{equation}
From this last equation we see that the propagator is
\begin{equation}
\label{ep2.32}
K_m(z,\vec{x}-\vec{x}^{'})=
\frac {1} {(2\pi)^{\nu}}
\left(\frac {z} {\epsilon}\right)^{\frac {\nu} {2}} 
\int \frac {{\cal K}_\mu(kz)}  {{\cal K}_\mu(k\epsilon)}
e^{-i\vec{k}\cdot (\vec{x}-\vec{x}^{'})}d^{\nu}k.
\end{equation}
As a consequence we can write
\begin{equation}
\label{ep2.33}
\phi(z,\vec{x})=
\int K_m(z,\vec{x}-{\vec{x}}^{'})
{\phi}_{\epsilon}({\vec{x}}^{'})d^{\nu}x^{'}.
\end{equation}
From (\ref{ep2.33}) we immediately gather that 
\begin{equation}
\label{ep2.34}
K_m(\epsilon,\vec{x}-{\vec{x}}^{'})=
\delta(\vec{x}-{\vec{x}}^{'}).
\end{equation}

\subsection{Approximate Massive Field Propagator}
We are now going to discuss a non-valid approach for the function
${\cal K}(k\epsilon)$. The issue here is that,  although $\epsilon$ is infinitesimal,
it can not  adopt the  $0-$value. As $k$ is an unbounded variable,
when $k\rightarrow\infty$, we have $k\epsilon\rightarrow\infty$. 
Notice first that
\begin{equation}
\label{ep2.35}
{\cal K}_\mu(k\epsilon)=\frac {2^{\mu-1}\Gamma(\mu)}
{(k\epsilon)^{\mu}}+O((k\epsilon)^{2-\mu}).
\end{equation}
We now make the approximation
\begin{equation}
\label{ep2.36}
{\cal K}_\mu(k\epsilon)=\frac {2^{\mu-1}\Gamma(\mu)}
{(k\epsilon)^{\mu}}.
\end{equation}
From (\ref{ep2.32}) we obtain an approximation for the propagator 
$K$ that we  shall call $M$.
We have then
\begin{equation}
\label{ep2.37}
M_m(z,\vec{x}-\vec{x}^{'})=
\frac {1} {(2\pi)^{\nu}}
\left(\frac {z} {\epsilon}\right)^{\frac {\nu} {2}} 
\frac {\epsilon^{\mu}} {2^{\mu-1}\Gamma(\mu)}
\int k^\mu {\cal K}_\mu(kz)
e^{-i\vec{k}\cdot (\vec{x}-\vec{x}^{'})}d^{\nu}k.
\end{equation}
Using again the Bochner formula we arrive at 
\begin{equation}
\label{ep2.38}
\int  k^\mu {\cal K}_\mu(kz)
e^{-i\vec{k}\cdot (\vec{x}-{\vec{x}}^{'})}d^{\nu}k=
\frac {(2\pi)^{\frac {\nu} {2}}} {|\vec{x}-\vec{x}^{'}|^{\frac {\nu} {2}-1}} 
\int\limits_0^{\infty} k^{\mu+\frac {\nu} {2}} 
{\cal K}_\mu(kz)
{\cal J}_{\frac {\nu} {2}-1}(k|\vec{x}-\vec{x}^{'}|)dk.
\end{equation}
By recourse to (\ref{ep2.19}) we then have
\begin{equation}
\label{ep2.39}
M_m(z,\vec{x}-\vec{x}^{'})=
\frac {{\epsilon}^{\mu-\frac {\nu} {2}}} {\pi^{\frac {\nu} {2}}}
\frac {\Gamma\left(\mu+\frac {\nu} {2}\right)}
{\Gamma(\mu)}
\left[\frac {z} {z^2+(\vec{x}-\vec{x}^{'})^2}\right]^{\mu+\frac {\nu} {2}}.
\end{equation}
We now define
\begin{equation}
\label{ep2.40}
\gamma=\frac {\nu} {2}+\mu=
\frac {\nu} {2}+\sqrt{\frac {\nu^2} {4}+\Delta(\Delta-\nu)},
\end{equation}
so that we can write
\begin{equation}
\label{ep2.41}
M_m(z,\vec{x}-\vec{x}^{'})=
\frac {{\epsilon}^{\gamma-\nu}} {\pi^{\frac {\nu} {2}}}
\frac {\Gamma\left(\gamma\right)}
{\Gamma\left(\gamma-\frac {\nu} {2}\right)}
\left[\frac {z} {z^2+(\vec{x}-\vec{x}^{'})^2}\right]^{\gamma}.
\end{equation}
We now realize that, by construction, 
\begin{equation}
\label{ep2.42}
M_m(\epsilon,\vec{x}-{\vec{x}}^{'})\neq
\delta(\vec{x}-{\vec{x}}^{'}),
\end{equation}
and define 
\begin{equation}
\label{ep2.43}
N_m(z,\vec{x}-\vec{x}^{'})=M_m(z,\vec{x}-\vec{x}^{'}){\epsilon}^{\nu-\gamma},
\end{equation}
which allows us to write  for $N_m$ the expression
\begin{equation}
\label{ep2.44}
N_m(z,\vec{x}-\vec{x}^{'})=
\frac {1} {\pi^{\frac {\nu} {2}}}
\frac {\Gamma\left(\gamma\right)}
{\Gamma\left(\gamma-\frac {\nu} {2}\right)}
\left[\frac {z} {z^2+(\vec{x}-\vec{x}^{'})^2}\right]^{\gamma}.
\end{equation}
Therefore, we have constructively proved that
\begin{equation}
\label{ep2.45}
\lim_{z\rightarrow 0} N_m(z,\vec{x}-{\vec{x}}^{'})\neq
\delta(\vec{x}-{\vec{x}}^{'}).
\end{equation}
Note that (\ref{ep2.44})  
 is indeed  the well known expression for the boundary-bulk
propagator for a scalar field in  configuration space.
However, this expression can only be used as an approximation
to the propagator $K$ when $\mu\cong\frac {\nu} {2}$.
   
\section{Minkowskian Case}

\setcounter{equation}{0}

\subsection{Massless Field Propagator}

Let is now deal with the case in which the boundary of the $ADS_{\nu+1}$ 
is the $\nu$-dimensional Minkowskian space. In the massless case
the field-equation  is

\begin{equation}
\label{ep3.1}
z^2\partial_z^2\hat{\phi}(z,k)+(1-\nu)z\partial_z\hat{\phi}(z,k)+z^2k^2\hat{\phi}(z,k)=0,
\end{equation}
where $k^2=k_0^2-{\vec{k}}^2=\rho$. Thus, we can write
\begin{equation}
\label{ep3.2}
z^2\partial_z^2\hat{\phi}(z,\rho)+(1-\nu)z\partial_z\hat{\phi}(z,\rho)+z^2\rho\hat{\phi}(z,\rho)=0,
\end{equation}
or, rewriting this last equation,
\begin{equation}
\label{ep3.3}
z^2\partial_z^2\hat{\phi}(z,\rho)+(1-\nu)z\partial_z\hat{\phi}(z,\rho)-
z^2\left[\mp i(\rho\pm i0)^{\frac {1} {2}}\right]^2\hat{\phi}(z,\rho)=0.
\end{equation}
The distribution  $(\rho\pm i0)^{\lambda}$ is defined as (see reference \cite{tp7})
\begin{equation}
\label{ep3.4}
(\rho\pm i0)^{\lambda}=\rho_+^{\lambda}+
e^{\pm i\pi\lambda}\rho_-^{\lambda},
\end{equation}
and can be cast in terms of  $H(x)$, the  Heaviside step function \cite{tp7}.
We recast now (\ref{ep3.3}) in the form of a Bessel equation
\begin{equation}
\label{ep3.5}
\partial_z^2\hat{\phi}(z,\rho)+\frac {(1-\nu)} {z}\partial_z\hat{\phi}(z,\rho)-
\left[\mp i(\rho\pm i0)^{\frac {1} {2}}\right]^2\hat{\phi}(z,\rho)=0.
\end{equation}
The solution of this equation that is i) regular at the origin and 2) vanishes for
$\rho\rightarrow\infty$,  becomes
\begin{equation}
\label{ep3.6}
\hat{\phi}(z,k)=
z^{\frac {\nu} {2}}{\cal K}_{\frac {\nu} {2}}[\mp i(\rho\pm i0)^{\frac {1} {2}}z].
\end{equation}
One  must  take into account that $\lim_{k\rightarrow\infty}e^{ikx}=0$ (see below 
in this section and \cite{jones}).
\begin{equation}
\label{ep3.7}
{\cal K}_{\frac {\nu} {2}}[\mp i(\rho\pm i0)^{\frac {1} {2}}z]=
\frac {2^{\frac {\nu} {2}-1}\Gamma\left(\frac {\nu} {2}\right)}
{[\mp i(\rho\pm i0)^{\frac {1} {2}}z]^{\frac {\nu} {2}}}+
O\left([\mp i(\rho\pm i0)^{\frac {1} {2}}z]^{-\frac {\nu} {2}+2}\right).
\end{equation}
We have then
\[\phi_\mp(z,x)=\frac {z^{\frac {\nu} {2}}} {(2\pi)^{\nu}}
\int a(\vec{k}) {\cal K}_{\frac {\nu} {2}}[\mp i(\rho\pm i0)^{\frac {1} {2}}z]
e^{-ik\cdot x}d^{\nu}k=\]
\begin{equation}
\label{ep3.8}
\int\hat{\phi}(z,k)e^{ik\cdot x}d^{\nu}k.
\end{equation}
From this last equation we deduce that
\begin{equation}
\label{ep3.9}
\phi_\mp(z,x)=\frac {z^{\frac {\nu} {2}}2^{1-\frac {\nu} {2}}} 
{(2\pi)^{\nu}\Gamma{\left(\frac {\nu} {2}\right)}}
\int [\mp i(\rho\pm i0)^{\frac {1} {2}}]^{\frac {\nu} {2}}
{\cal K}_{\frac {\nu} {2}}[\mp i(\rho\pm i0)^{\frac {1} {2}}z]
\hat{\phi}_0(k)
e^{-ik\cdot x}d^{\nu}k,
\end{equation}
or, equivalently,
\[\phi_\mp(z,x)=\frac {z^{\frac {\nu} {2}}2^{1-\frac {\nu} {2}}} 
{(2\pi)^{\nu}\Gamma{\left(\frac {\nu} {2}\right)}}
\int\int [\mp i(\rho\pm i0)^{\frac {1} {2}}]^{\frac {\nu} {2}}
{\cal K}_{\frac {\nu} {2}}[\mp i(\rho\pm i0)^{\frac {1} {2}}z]\times\]
\begin{equation}
\label{ep3.10}
{\phi}_0(x^{'})
e^{-ik\cdot (x-x^{'})}d^{\nu}kd^{\nu}x^{'}.
\end{equation}
The ensuing propagator becomes then 
\begin{equation}
\label{ep3.11}
K_\mp(z,x-x^{'})=\frac {z^{\frac {\nu} {2}}2^{1-\frac {\nu} {2}}} 
{(2\pi)^{\nu}\Gamma{\left(\frac {\nu} {2}\right)}}
\int [\mp i(\rho\pm i0)^{\frac {1} {2}}]^{\frac {\nu} {2}}
{\cal K}_{\frac {\nu} {2}}[\mp i(\rho\pm i0)^{\frac {1} {2}}z]
e^{-i\vec{k}\cdot (x-x^{'})}d^{\nu}k.
\end{equation}
Thus, the corresponding Feynman's propagator is
\begin{equation}
\label{ep3.12}
K_{F}(z,x-x^{'})=\frac {z^{\frac {\nu} {2}}2^{1-\frac {\nu} {2}}} 
{(2\pi)^{\nu}\Gamma{\left(\frac {\nu} {2}\right)}}
\int [-i(\rho+i0)^{\frac {1} {2}}]^{\frac {\nu} {2}}
{\cal K}_{\frac {\nu} {2}}[-i(\rho+i0)^{\frac {1} {2}}z]
e^{-i\vec{k}\cdot (x-x^{'})}d^{\nu}k.
\end{equation}
Note that the Feynman propagator is a function of $\rho+i0$, as it should.
For the anti-Feynman propagator we have instead
\begin{equation}
\label{ep3.13}
K_{AF}(z,x-x^{'})=\frac {z^{\frac {\nu} {2}}2^{1-\frac {\nu} {2}}} 
{(2\pi)^{\nu}\Gamma{\left(\frac {\nu} {2}\right)}}
\int [i(\rho-i0)^{\frac {1} {2}}]^{\frac {\nu} {2}}
{\cal K}_{\frac {\nu} {2}}[i(\rho-i0)^{\frac {1} {2}}z]
e^{-i\vec{k}\cdot (x-x^{'})}d^{\nu}k.
\end{equation}
The expression for the Wheeler's propagator is:
\begin{equation}
\label{ep3.14}
W(z,x-x^{'})=
\frac {1} {2}[K_{F}(z,x-x^{'})+K_{AF}(z,x-x^{'})].
\end{equation}
Using the relations
\begin{equation}
\label{ep3.15}
K_F(z,\rho)=\frac {z^{\frac {\nu} {2}}2^{1-\frac {\nu} {2}}} 
{\Gamma{\left(\frac {\nu} {2}\right)}}
[-i(\rho+i0)^{\frac {1} {2}}]^{\frac {\nu} {2}}
{\cal K}_{\frac {\nu} {2}}[-i(\rho+i0)^{\frac {1} {2}}z],
\end{equation}
and
\begin{equation}
\label{ep3.16}
K_{AF}(z,\rho)=\frac {z^{\frac {\nu} {2}}2^{1-\frac {\nu} {2}}} 
{\Gamma{\left(\frac {\nu} {2}\right)}}
[i(\rho-i0)^{\frac {1} {2}}]^{\frac {\nu} {2}}
{\cal K}_{\frac {\nu} {2}}[i(\rho-i0)^{\frac {1} {2}}z],
\end{equation}
we can define, as  usual, the retarded propagator
\begin{equation}
\label{ep3.17}
K_R(z,\rho)=H(k^0)
K_{F}(z,\rho)+H(-k^0)K_{AF}(z,\rho),
\end{equation}
and the advanced propagator
\begin{equation}
\label{ep3.18}
K_A(z,\rho)=H(k^0)
K_{AF}(z,\rho)+H(-k^0)K_{F}(z,\rho).
\end{equation}
We are going to show now that $\lim\limits_{k\rightarrow\infty}e^{ikx}=0$ (see \cite{jones})
Let $\hat{\phi}$ be a test function  belonging to a sub-space  ${\cal S}$
of Schwartz's one \cite{tp4,tp7}. Its Fourier transform is
\begin{equation}
\label{ep3.19}
\phi(k)=\int\limits_{-\infty}^{\infty}
\hat{\phi(x)}e^{ikx}dx,
\end{equation}
where $\phi$ belongs to ${\cal S}$. Then one can  verify that
\begin{equation}
\label{ep3.20}
0=\lim\limits_{k\rightarrow\infty}\phi(k)=
\lim\limits_{k\rightarrow\infty}\int\limits_{-\infty}^{\infty}
\hat{\phi(x)}e^{ikx}dx=
\int\limits_{-\infty}^{\infty}
\hat{\phi(x)}\lim\limits_{k\rightarrow\infty}e^{ikx}dx.
\end{equation}
As a consequence, we obtain
\begin{equation}
\label{ep3.21}
\lim\limits_{k\rightarrow\infty}e^{ikx}=0
\end{equation}
The Feynman propagator is, according to (\ref{ep3.12},)
\begin{equation}
\label{ep3.22}
K_F(z,x)=\frac {z^{\frac {\nu} {2}}2^{1-\frac {\nu} {2}}} 
{(2\pi)^\nu\Gamma{\left(\frac {\nu} {2}\right)}}\int
[-i(\rho+i0)^{\frac {1} {2}}]^{\frac {\nu} {2}}
{\cal K}_{\frac {\nu} {2}}[-i(\rho+i0)^{\frac {1} {2}}z]e^{-ik\cdot x}d^\nu k.
\end{equation}
Since ${\cal K}_{\frac {\nu} {2}}$ is exponentially decreasing or oscillating, we can evaluate the integral that defines $ K_F $ by means of a Wick rotation over $k_0$. Therefore we have the change of variables
$k_0=ik_ {0E}$, $x_0=ix_{0E}$, $k_E^2 = k_{0E}^2+\vec {k}^2$, and 
$x_E^2=x_{0E}^2+\vec {x}^2$. Casting the integral that defines the propagator in terms of these new 
variables, we obtain
\begin{equation}
\label{ep3.23}
K_F(z,\vec{x}_E)=\frac {iz^{\frac {\nu} {2}}2^{1-\frac {\nu} {2}}} 
{(2\pi)^\nu\Gamma{\left(\frac {\nu} {2}\right)}}\int
k_E^{\frac {\nu} {2}}
{\cal K}_{\frac {\nu} {2}}(k_Ez)e^{-i\vec{k}_E\cdot \vec{x}_E}d^\nu k_E.
\end{equation}
Using Bochner's formula together with  (\ref{ep2.19}) we have 
\begin{equation}
\label{ep3.24}
K_F(z,x_E)=\frac {i\Gamma(\nu)} {\pi^{\frac {\nu} {2}}
\Gamma\left(\frac {\nu} {2}\right)}
\left[\frac {z} {z^2+x_E^2}\right]^\nu.
\end{equation}
Now, making the change to Minkowskian variables and taking into account that the Fourier
 transform of a distribution that depends on $\rho-i0$ is a distribution that 
 depends on $x^2+i0$, 
  we obtain
\begin{equation}
\label{ep3.25}
K_F(z,x)=\frac {i\Gamma(\nu)} {\pi^{\frac {\nu} {2}}
\Gamma\left(\frac {\nu} {2}\right)}
\left[\frac {z} {z^2-x^2-i0}\right]^\nu,
\end{equation}
which is the expression of the Feynman propagator in terms of the
variables of the configuration space. For the 
anti-Feynman propagator we analogously find
\begin{equation}
\label{ep3.26}
K_{AF}(z,x)=\frac {i\Gamma(\nu)} {\pi^{\frac {\nu} {2}}
\Gamma\left(\frac {\nu} {2}\right)}
\left[\frac {z} {z^2-x^2+i0}\right]^\nu.
\end{equation}

\subsection{Massive Field Propagator}

For the massive case, the field-motion equation  is
\begin{equation}
\label{ep3.27}
\partial_z^2\hat{\phi}(z,\rho)+\frac {(1-\nu)} {z}\partial_z\hat{\phi}(z,\rho)-
\left\{\left[\mp i(\rho\pm i0)^{\frac {1} {2}}\right]^2+
\frac {\Delta(\Delta-\nu)} {z^2}\right\}\hat{\phi}(z,\rho)=0,\end{equation}
with, again,
\begin{equation}
\label{ep3.28}
\mu=\sqrt{\frac {\nu^2} {4}+\Delta(\Delta-\nu)}.
\end{equation}
The pertinent solution is now
\begin{equation}
\label{ep3.29}
{\hat{\phi}}_\mp(z,\rho)=z^{\frac {\nu} {2}}{\cal K}_{\mu}[\mp i(\rho\pm i0)^{\frac {1} {2}}z].
\end{equation}
The field-expression  in  configuration space is then
\begin{equation}
\label{ep3.30}
\phi_\mp(z,x)=\frac {z^{\frac {\nu} {2}}} {(2\pi)^{\nu}}
\int a(\vec{k}) {\cal K}_{\mu}[\mp i(\rho\pm i0)^{\frac {1} {2}}z]
e^{-ik\cdot x}d^{\nu}k.
\end{equation}
Once again we choose
\[\phi(\epsilon,x)=\phi_{\epsilon}(x)=
\frac {{\epsilon}^{\frac {\nu} {2}}} {(2\pi)^{\nu}}
\int a(\vec{k})  {\cal K}_{\mu}[\mp i(\rho\pm i0)^{\frac {1} {2}}\epsilon]
e^{-ik\cdot x}d^{\nu}k=\]
\begin{equation}
\label{ep3.31}
\frac {1} {(2\pi)^\nu}
\int \hat{\phi}_{\epsilon}(k)e^{-ik\cdot x}d^{\nu}k,
\end{equation}
and from (\ref{ep3.23}) we obtain
\begin{equation}
\label{ep3.32}
a(k)=\frac {\hat{\phi}_{\epsilon}(k)}
{{\epsilon}^{\frac {\nu} {2}}{\cal K}_{\mu}[\mp i(\rho\pm i0)^{\frac {1} {2}}\epsilon]}.
\end{equation}
We have then the following relation for the solution
\begin{equation}
\label{ep3.33}
\phi_\mp(z,x)=
\frac {1} {(2\pi)^{\nu}}
\left(\frac {z} {\epsilon}\right)^{\frac {\nu} {2}} 
\int\int \frac {{\cal K}_{\mu}[\mp i(\rho\pm i0)^{\frac {1} {2}}z]}
{{\cal K}_{\mu}[\mp i(\rho\pm i0)^{\frac {1} {2}}\epsilon]}
\phi_{\epsilon}(x^{'})
e^{-ik\cdot (x-x^{'})}d^{\nu}kd^{\nu}x^{'},
\end{equation}
so that the propagator is now
\begin{equation}
\label{ep3.34}
K_{m\mp}(z,x-x^{'})=
\frac {1} {(2\pi)^{\nu}}
\left(\frac {z} {\epsilon}\right)^{\frac {\nu} {2}} 
\int \frac {{\cal K}_{\mu}[\mp i(\rho\pm i0)^{\frac {1} {2}}z]}
{{\cal K}_{\mu}[\mp i(\rho\pm i0)^{\frac {1} {2}}\epsilon]}
e^{-i\vec{k}\cdot (x-x^{'})}d^{\nu}k.
\end{equation}
The corresponding Feynman's propagator becomes
\begin{equation}
\label{ep3.35}
K_{mF}(z,x-x^{'})=
\frac {1} {(2\pi)^{\nu}}
\left(\frac {z} {\epsilon}\right)^{\frac {\nu} {2}} 
\int \frac {{\cal K}_{\mu}[-i(\rho+i0)^{\frac {1} {2}}z]}
{{\cal K}_{\mu}[-i(\rho+i0)^{\frac {1} {2}}\epsilon]}
e^{-i\vec{k}\cdot (x-x^{'})}d^{\nu}k,
\end{equation}
and for the anti-Feynman propagator we obtain the expression
\begin{equation}
\label{ep3.36}
K_{mAF}(z,x-x^{'})=
\frac {1} {(2\pi)^{\nu}}
\left(\frac {z} {\epsilon}\right)^{\frac {\nu} {2}} 
\int \frac {{\cal K}_{\mu}[i(\rho-i0)^{\frac {1} {2}}z]}
{{\cal K}_{\mu}[ i(\rho-i0)^{\frac {1} {2}}\epsilon]}
e^{-i\vec{k}\cdot (x-x^{'})}d^{\nu}k.
\end{equation}
Finally, the definition of Wheeler propagators, retarded and advanced, 
is similar to that of the preceding subsection.

\subsection{Approximation}
We now evaluate in approximate fashion the  propagator 
${{\cal K}_{\mu}[-i(\rho+i0)^{\frac {1} {2}}\epsilon]}$

\begin{equation}
\label{ep3.37}
{{\cal K}_{\mu}[-i(\rho+i0)^{\frac {1} {2}}\epsilon]}=
\frac {2^{\mu-1}\Gamma(\mu)}
{(-i)^\mu(\rho+i0)^{\frac {\mu} {2}}\epsilon^\mu},
\end{equation}
entailing
\begin{equation}
\label{ep3.38}
M_{mF}(z,x)=
\frac {z^{\frac {\nu} {2}}} {(2\pi)^{\nu}}
\frac {\epsilon^{\mu-\frac {\nu} {2}}} 
{2^{\mu-1}\Gamma(\mu)}
\int {\cal K}_{\mu}[-i(\rho+i0)^{\frac {1} {2}}z]
[-i(\rho+i0)^{\frac {1} {2}}]^{\mu}
e^{-ik\cdot x}d^{\nu}k.
\end{equation}
Effecting again the above Wick's rotation we obtain
\begin{equation}
\label{ep3.39}
M_{mF}(z,\vec{x}_E)=
\frac {iz^{\frac {\nu} {2}}} {(2\pi)^{\nu}}
\frac {\epsilon^{\mu-\frac {\nu} {2}}} 
{2^{\mu-1}\Gamma(\mu)}
\int k_E^\mu{\cal K}_{\mu}(k_Ez)
e^{-i\vec{k}_E\cdot \vec{x}_E}d^{\nu}k_E.
\end{equation}
This integral is evaluated as  in the previous cases. One has

\begin{equation}
\label{ep3.40}
M_{mF}(z,\vec{x}_E)=
\frac {i\epsilon^{\mu-\frac {\nu} {2}}} {\pi^{\frac {\nu} {2}}}
\frac {\Gamma\left(\mu-\frac {\nu} {2}\right)}
{\Gamma(\mu)}
\left(\frac {z} {z^2+x_E^2}\right)^{\mu+\frac {\nu} {2}}.
\end{equation}
Changing  variables as above we arrive at 
\begin{equation}
\label{ep3.41}
M_{mF}(z,x)=
\frac {i\epsilon^{\gamma-\nu}} {\pi^{\frac {\nu} {2}}}
\frac {\Gamma\left(\gamma\right)}
{\Gamma\left(\gamma-\frac {\nu} {2}\right)}
\left(\frac {z} {z^2-x^2-i0}\right)^{\gamma},
\end{equation}
where
\begin{equation}
\label{eq3.42}
\gamma=\frac {\nu} {2}+\sqrt{\frac {\nu^2} {4}+\Delta(\Delta-\nu)}.
\end{equation}
Now we return to  the inequality
\begin{equation}
\label{ep3.43}
M_{mF}(\epsilon,x)\neq\delta(x)
\end{equation}
The following relation  is valid for $N_F$
\begin{equation}
\label{ep3.44}
N_{mF}(\epsilon,x)=\epsilon^{\nu-\gamma}M_{mF}(\epsilon,x).
\end{equation}
Proceeding in analogous fashion with the Anti-Feynman propagator we obtain the approximation
\begin{equation}
\label{ep3.45}
M_{mAF}(z,x)=
\frac {i\epsilon^{\gamma-\nu}} {\pi^{\frac {\nu} {2}}}
\frac {\Gamma\left(\gamma\right)}
{\Gamma\left(\gamma-\frac {\nu} {2}\right)}
\left(\frac {z} {z^2-x^2+i0}\right)^{\gamma}
\end{equation}

\setcounter{equation}{0}

\section{Two points correlation functions in Euclidean space}

\subsection{Massless case}

To evaluate the two-points correlation function of a scalar operator
we use the result obtained in \cite{nas}. This is
\begin{equation}
\label{eq4.1}
<{\cal O}(\vec{x}_1){\cal O}(\vec{x}_2)>=-
\int \sqrt{g}\partial_\mu K(y_0,\vec{y}-\vec{x}_1)
\partial^\mu K(y_0,\vec{y}-\vec{x}_2)d^{\nu+1} y,
\end{equation}
where $0\leq y_0=z<\infty$, $y_\mu=x_\mu$, $\mu\neq 0$, and then
\begin{equation}
\label{eq4.2}
<{\cal O}(\vec{x}_1){\cal O}(\vec{x}_2)>=-
\int\limits_{Boundary}\lim_{z\rightarrow 0}[
z^{1-\nu}
K(z,\vec{x}-\vec{x}_1)
\partial_z K(z,\vec{x}-\vec{x}_2)]d^\nu x.
\end{equation}
As $\lim\limits_{z\rightarrow 0} K(0,\vec{x}_1-\vec{x}_2)=\delta(\vec{x}_1-\vec{x}_2)$,
we obtain
\begin{equation}
\label{eq4.3}
<{\cal O}(\vec{x}_1){\cal O}(\vec{x}_2)>=-
\lim_{z\rightarrow 0}[
z^{1-\nu}
\partial_z K(z,\vec{x}_1-\vec{x}_2)].
\end{equation}
Using now the expression for $K$ given in Eq.(\ref{ep2.20}) we have
\begin{equation}
\label{eq4.4}
<{\cal O}(\vec{x}_1){\cal O}(\vec{x}_2)>=-
\frac {\Gamma(\nu+1)} {\pi^{\frac {\nu} {2}}\Gamma\left(
\frac {\nu} {2}\right)}\frac {1} {(\vec{x}_1-\vec{x}_2)^{2\nu}}.
\end{equation}
Accordingly,  we have here arrived to the usual, well-known result.

\subsection{Massive case}

For the massive case we obtain similarly
\begin{equation}
\label{eq4.5}
<{\cal O}(\vec{x}_1){\cal O}(\vec{x}_2)>_m=-
\int\limits_{Boundary}
[z^{1-\nu}
K_m(z,\vec{x}-\vec{x}_1)
\partial_z K(z,\vec{x}-\vec{x}_2)]d^\nu x.
\end{equation}
As $K_m(\epsilon,\vec{x}-\vec{x}_1)=\delta(\vec{x}-\vec{x}_1)$
we can write
\begin{equation}
\label{eq4.6}
<{\cal O}(\vec{x}_1){\cal O}(\vec{x}_2)>_m=-\int
\delta(\vec{x}-\vec{x}_1)
[z^{1-\nu}
\partial_z K(z,\vec{x}-\vec{x}_2)]_{z=\epsilon}d^\nu x.
\end{equation}
Thus we arrive at
\begin{equation}
\label{eq4.7}
<{\cal O}(\vec{x}_1){\cal O}(\vec{x}_2)>_m=-
[z^{1-\nu}
\partial_z K_m(z,\vec{x}_1-\vec{x}_2)]_{z=\epsilon}.
\end{equation}
Now, we use the expression for $K_m$ given in (\ref{ep2.32}) and write
\begin{equation}
\label{eq4.8}
<{\cal O}(\vec{x}_1){\cal O}(\vec{x}_2)>_m=-\frac {\epsilon^{1-\nu}}
{(2\pi)^{\nu}}\partial_z\left[\left(\frac {z} {\epsilon}\right)^{\frac {\nu} {2}}
\int\frac {{\cal K}_\mu(kz)} {{\cal K}_\mu(k\epsilon)}
e^{-i\vec{k}\cdot(\vec{x}_1-\vec{x}_2)}d^\nu k \right]_{z=\epsilon},
\end{equation}
or, equivalently,
\[<{\cal O}(\vec{x}_1){\cal O}(\vec{x}_2)>_m=-\frac {\epsilon^{1-\frac {3\nu} {2}}}
{(2\pi)^{\nu}}\left[\frac {\nu} {2}z^{\frac {\nu} {2}-1}
\int\frac {{\cal K}_\mu(kz)} {{\cal K}_\mu(k\epsilon)}
e^{-i\vec{k}\cdot(\vec{x}_1-\vec{x}_2)}d^\nu k \right.+\]
\begin{equation}
\label{eq4.9}
\left.z^{\frac {\nu} {2}}
\int k\frac {{\cal K}^{'}_\mu(kz)} {{\cal K}_\mu(k\epsilon)}
e^{-i\vec{k}\cdot(\vec{x}_1-\vec{x}_2)}d^\nu k \right]_{z=\epsilon}.
\end{equation}
Using now the following result, given in \cite{tp13},
\begin{equation}
\label{eq4.10}
{\cal K}^{'}_\mu(z)=-\frac {\mu} {z}{\cal K}_\mu+{\cal K}_{\mu-1},
\end{equation}
we obtain
\begin{equation}
\label{eq4.11}
<{\cal O}(\vec{x}_1){\cal O}(\vec{x}_2)>=\epsilon^{-\nu}(\mu-\frac {\nu} {2})
\delta(\vec{x}_1-\vec{x}_2)-
\frac {\epsilon^{1-\nu}} {(2\pi)^\nu}
\int k\frac {{\cal K}_{\mu-1}(k\epsilon)} {{\cal K}_\mu(k\epsilon)}
e^{-i\vec{k}\cdot(\vec{x}_1-\vec{x}_2)}d^\nu k.
\end{equation}

Note that we have not renormalized the correlation functions. We will do that using the results of 
\cite{z2} in a forthcoming  paper.

\setcounter{equation}{0}

\section{Two points correlation functions in Minkowskian space}

\subsection{Massless case}

Similarly to the Euclidean case we obtain for the Minkowskian one the result
\begin{equation}
\label{eq5.1}
<{\cal O}(x_1){\cal O}(x_2)>_F=i
\lim_{z\rightarrow 0}[
z^{1-\nu}
\partial_z K_F(z,x_1-x_2)].
\end{equation}
Thus, we obtain for the Feynman's propagator
\begin{equation}
\label{eq5.2}
<{\cal O}(\vec{x}_1){\cal O}(\vec{x}_2)>_F=-
\frac {\Gamma(\nu+1)} {\pi^{\frac {\nu} {2}}\Gamma\left(
\frac {\nu} {2}\right)}\frac {1} {[(\vec{x}_1-\vec{x}_2)^2-
(x_{10}-x_{20})^2+i0]^\nu},
\end{equation}
and for Anti-Feynman one
\begin{equation}
\label{eq5.3}
<{\cal O}(\vec{x}_1){\cal O}(\vec{x}_2)>_{AF}=-
\frac {\Gamma(\nu+1)} {\pi^{\frac {\nu} {2}}\Gamma\left(
\frac {\nu} {2}\right)}\frac {1} {[(\vec{x}_1-\vec{x}_2)^2-
(x_{10}-x_{20})^2-i0]^\nu}.
\end{equation}

\subsection{Massive case}

Again, following the developments of the Euclidean case, we have, for the Minkowskian instance,
 the two points Feynman's correlator 
\begin{equation}
\label{eq5.4}
<{\cal O}(\vec{x}_1){\cal O}(\vec{x}_2)>_{mF}=i
[z^{1-\nu}
\partial_z K_{mF}(z,x_1-x_2)]_{z=\epsilon}.
\end{equation}
Thus, we have
\begin{equation}
\label{eq5.5}
<{\cal O}(x_1){\cal O}(x_2)>_{mF}=i\frac {\epsilon^{1-\nu}}
{(2\pi)^{\nu}}\partial_z\left[\left(\frac {z} {\epsilon}\right)^{\frac {\nu} {2}}
\int\frac {{\cal K}_\mu[-i(\rho+i0)^{\frac {1} {2}}z]} 
{{\cal K}_\mu[-i(\rho+i0)^{\frac {1} {2}}\epsilon]}
e^{-ik\cdot(x_1-x_2)}d^\nu k \right]_{z=\epsilon},
\end{equation}
or equivalently,
\[<{\cal O}(x_1){\cal O}(x_2)>_{mF}=i\frac {\epsilon^{1-\frac {3\nu} {2}}}
{(2\pi)^{\nu}}\left[\frac {\nu} {2}z^{\frac {\nu} {2}-1}
\int\frac {{\cal K}_\mu[-i(\rho+i0)^{\frac {1} {2}}z]} 
{{\cal K}_\mu[-i(\rho+i0)^{\frac {1} {2}}\epsilon]}
e^{-ik\cdot(x_1-x_2)}d^\nu k \right.-\]
\begin{equation}
\label{eq5.6}
\left.iz^{\frac {\nu} {2}}
\int (\rho+i0)^{\frac {1} {2}}\frac {{\cal K}^{'}_\mu[-i(\rho+i0)^{\frac {1} {2}}z]} 
{{\cal K}_\mu[-i(\rho-i0)^{\frac {1} {2}}\epsilon]}
e^{-ik\cdot(x_1-x_2)}d^\nu k \right]_{z=\epsilon}.
\end{equation}
Using again (\ref{eq4.10}) we finally obtain 
\[<{\cal O}(x_1){\cal O}(x_2)>_F=\epsilon^{-\nu}(\mu-\frac {\nu} {2})
\delta(x_1-x_2)+\]
\begin{equation}
\label{eq5.7}
\frac {\epsilon^{1-\nu}} {(2\pi)^\nu}
\int (\rho+i0)^{\frac {1} {2}}\frac {{\cal K}_{\mu-1}[-i(\rho+i0)^{\frac {1} {2}}\epsilon]} 
{{\cal K}_\mu[-i(\rho+i0)^{\frac {1} {2}}\epsilon]}
e^{-ik\cdot(x_1-x_2)}d^\nu k.
\end{equation}
For the Anti-Feynman propagator we obtain in  analogous fashion
\[<{\cal O}(x_1){\cal O}(x_2)>_{mAF}=\epsilon^{-\nu}(\mu-\frac {\nu} {2})
\delta(x_1-x_2)-\]
\begin{equation}
\label{eq5.8}
\frac {\epsilon^{1-\nu}} {(2\pi)^\nu}
\int (\rho-i0)^{\frac {1} {2}}\frac {{\cal K}_{\mu-1}[i(\rho-i0)^{\frac {1} {2}}\epsilon]} 
{{\cal K}_\mu[i(\rho-i0)^{\frac {1} {2}}\epsilon]}
e^{-ik\cdot(x_1-x_2)}d^\nu k.
\end{equation}

\nd Note that we have not renormalized the correlation functions here. 
We will do so using the results of \cite{z2} in a forthcoming  paper.
We strongly emphasize the fact that holographic renormalization should not be used in this context because 
  it constitutes a non-rigorous mathematical technique.

\section{Discussion}

\nd
In this work we have first calculated, without using any conjecture, the 
boundary-bulk Feynman, Anti-Feynman, and Wheeler propagators for both  a 
massless and a massive scalar field, by recourse to the theory of distributions.

\nd
Note that  such a conjecture was made in  2002 by Son and Starinets \cite{son}, only for 
the Feynman propagator,  and that they did not get a correct result

\nd
As further novelties, in the paper we show that, for massive scalar fields,
the expression for the boundary-bulk propagator in  Euclidean momentum space does not correspond to 
the expression used in the configuration space, but it is rather  a mere approximation.

\nd
Subsequently, using the previous results, we have evaluated
the correlation functions of scalar operators
corresponding to massless and massive scalar fields.

\nd
{\bf Unlike the results obtained in \cite{son}, with the ones obtained
here you can calculate the n-points correlation functions from gravity. This is feasible  for a scalar 
operator when $n$ is an arbitrary natural number.}

\end{document}